OpenStudy (anonymous):

what is the derivative of arctan(x/3)

OpenStudy (anonymous):

do you use 1 / 1 + x^2

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Just like x/2

OpenStudy (anonymous):

the der. of arctan x = 1/ (1+ x^2)

OpenStudy (anonymous):

Using the chain rule, the der. of arctan x/2 is (1/2) * (1 / 1+ (x/2)^2)

OpenStudy (anonymous):

where did the (1/2) come from? the exp.-1

OpenStudy (anonymous):

find the derivative of: 1. y = sec^-1(t+1) 2. y = e^(cos^-1) 3. y = csc^-1(x^-1)

OpenStudy (anonymous):

for #1, do you use sec^-1x= 1 / I x I sq root x^2 -1 ( absolute value : I x I )